approximate the real root to four decimal places of x3+5x-3=0 (newton raphson method)


1
Expert's answer
2022-01-17T07:02:36-0500

Solution:

Here "x^{3}+5 x-3=0"

Let "f(x)=x^{3}+5 x-3"

"\\

\\\frac{d}{d x}\left(x^{3}+5 x-3\right)=3 x^{2}+5

\therefore f(x)=3 x^{2}+5"

Here


Here "f(0)=-3<0\ and\ f(1)=3>0"

"\therefore" Root lies between 0 and 1

"x_{0}=\frac{0+1}{2}=0.5

\\x_{0}=0.5"

"1^{s t}" iteration :

"f\left(x_{0}\right)=f(0.5)=0.5^{3}+5 \cdot 0.5-3=-0.375

\\f^{\prime}\left(x_{0}\right)=f(0.5)=3 \cdot 0.5^{2}+5=5.75"

"\begin{aligned}

&x_{1}=x_{0}-\frac{f\left(x_{0}\right)}{f\left(x_{0}\right)} \\

&x_{1}=0.5-\frac{-0.375}{5.75} \\

&x_{1}=0.5652

\end{aligned}"

"2^{n d}" iteration :

"\begin{aligned}

&f\left(x_{1}\right)=f(0.5652)=0.5652^{3}+5 \cdot 0.5652-3=0.0067 \\

&f^{\prime}\left(x_{1}\right)=f(0.5652)=3 \cdot 0.5652^{2}+5=5.9584

\end{aligned}"

"\begin{aligned}

&x_{2}=x_{1}-\frac{f\left(x_{1}\right)}{f\left(x_{1}\right)} \\

&x_{2}=0.5652-\frac{0.0067}{5.9584} \\

&x_{2}=0.5641

\end{aligned}"

"3^{\text {rd }}" iteration :

"f\left(x_{2}\right)=f(0.5641)=0.5641^{3}+5 \cdot 0.5641-3=0"

"\begin{aligned}
f^{\prime}\left(x_{2}\right)=f(0.5641)=3 \cdot 0.5641^{2}+5=5.9546 \\
x_{3}=x_{2}-\frac{f\left(x_{2}\right)}{f\left(x_{2}\right)} \\
x_{3}=0.5641-\frac{0}{5.9546} \\
x_{3}=0.5641 \\
\text { Approximate root of the equation } x^{3}+5 x-3=0 
\\\text { using Newton Raphson mehtod is } 0.5641 \text { (After } 3 \text { iterations) }
\end{aligned}"